# MAnycast Reloaded: a Tool for an Open, Fast, Responsible and Efficient Daily Anycast Census

### v1

@@ 5.9. Lessons learned
We find the performance of both anycast-based and GCD measurements depend on the number, and geographical and topological diversity of sites. Since the majority of anycast is from hypergiants with global deployments, we observe that few nodes on different continents can detect the majority of anycast, as also exemplified by BGPTools (BGPTools, 2025). However, for detecting regional anycast a larger measurement platform is required. Furthermore, enumeration and geolocation of anycast sites requires a well-distributed measurement platform.
@@
We find our census provides good coverage of the anycast landscape, as confirmed by extensive validation. However, for IPv6 we are restricted by hitlist coverage. Additionally, for enumeration we find our census provides a lower bound as latency-based measurements cannot differentiate between anycast sites in near-proximity.
@@
When using our census, we advise to keep the FP limitation of the anycast-based approach in mind, especially when only two VPs receive responses. As for GCD, we find it is highly accurate but has rare cases of FNs due to regional anycast.
@@
We encourage operators to run their own census using the MAnycastR tool, which supports both anycast- and latency-based measurements. Furthermore, our anycast-based target lists are made available in our public census, making it possible to combine them with other AT lists to improve coverage.
@@ 6. Conclusions and future work
In this work we presented MAnycast Reloaded (MAnycastR), a clean slate reimplementation inspired by the MAnycast2 anycast measurement approach of Sommese et al. (Sommese et al., 2020) and by the Great Circle Distance (GCD) approach pioneered by Cicalese et al. with iGreedy (Cicalese et al., 2015). Through extensive validation we show that MAnycastR allows for an efficient, accurate and precise daily anycast census for both IPv4 and IPv6 and with support for transport layer probing using UDP/DNS and TCP. We release this daily census to the community through a public Git repository1 and release the entire MAnycastR toolchain under a permissive open source licence [^4] LOCATION ANONYMISED FOR REVIEW.
We intend to further extend MAnycastR in future work. Specifically, we want to check responsiveness from a single VP before probing from all VPs, add support for a canary anycast deployment to detect outages, trigger-based detection of temporary anycast – e.g., from BGP route collectors, add GCD support using UDP and more. We are also planning to analyse longitudinal anycast dynamics and work in which we use MAnycastR to detect suspected BGP hijacking.
Appendix A GCD using RIPE Atlas
@@
We performed GCD using RIPE Atlas for the 23,821 ATs of 16 September, 2024. This measurement lasted three days due to rate-limiting, even with increased spending quotas. Furthermore, it incurred a cost of 37 million RIPE Atlas credits which would take a volunteer hosting a RIPE Atlas probe approximately five years to accumulate. Finally, results show inconsistency in detection and enumeration of anycast due to non-guaranteed availability of RIPE Atlas VPs. For these reasons we find RIPE Atlas unsuitable for the daily census.
@@
However, as discussed in §5.2 we do find that RIPE Atlas achieves better enumeration. However, this comes at a price in terms of probing cost. This can be seen in Figure 8, where we plot the number of PoPs detected for a Cloudflare prefix, with presence in 300+ cities, and the probing cost when decreasing the inter-node distance. Our RIPE Atlas scan was performed using 481 VPs with at least 100km inter-VP distances to maximize geographical coverage. By increasing the inter-VP distance up to 1,000km we reduce the number of VPs. We find that there is a linear increase in enumeration capabilities, whereas the probing cost shows an exponential increase.
@@
Appendix B Increasing GCD coverage
As mentioned, the GCDArk measurement used more VPs as it was performed in the development environment of Ark. The regular GCD measurement targeting the ATs uses 163 VPs, whereas the GCDArk measurement targeting the full hitlist used an additional 64 VPs totaling to 227. Figure 9 plots the number of sites detected behind anycast prefixes for the different number of VPs. We find that the additional VPs increases enumeration capabilities from $\sim$55 to $\sim$65, an increase of 18%. The probing cost, which is linear with the number of VPs, increases 39%. Unlike RIPE Atlas, we observe that results remain consistent and daily measuring of ATs remains feasible despite the increase in VPs.
As the Ark platform grows, we expect our census to increase performance in detection, enumeration, and geolocation of anycast using the GCD approach. Future work is to selectively use Ark VPs based on geographical coverage, similar to how we selected RIPE Atlas VPs, to reduce the probing cost of our measurement.
Figure 9. Plotting enumeration capabilities of the current MAnycastR GCD measurement and that with the development Ark nodes.
@@ C.1. Detection
As mentioned in §2.3, prior work attempted to detect and enumerate anycast using CHAOS queries. We build on this prior work and leverage CHAOS queries as an opportunity to perform a side-by-side comparison in isolation of the anycast-based approach in MAnycastR and the built-in GCD approach. We take our nameserver hitlist as a starting point. We send CHAOS queries to each target from each of our 32 VPs according to RFC 4892 (Woolf and Conrad, 2007) to the IPv4 and IPv6 addresses on the hitlist and record the responses. Then we run a separate synchronised anycast-based measurement with 1 second offsets against the hitlist as well as a separate GCD-based measurement. In other words: we run both the anycast-based and GCD measurement toward all addresses using the MAnycastR deployment. Of the 161 K nameservers probed, we detect 2,762 to be anycast using the anycast-based approach of which 2,371 are also found using GCD.
We also observe 414 nameservers that are unresponsive to ICMP and TCP, hence the GCD method cannot detect them. 50 of these prefixes are detected to be anycast using the anycast-based method, we observe they give strong indication of being TPs as they are captured at ¿ 3 VPs. These 50 include prefixes from well-known operators, such as eBay and Oracle, highlighting the utility of the added protocol support in MAnycastR.
@@ C.2. Enumeration
As mentioned, the MAnycastR tool has built-in support for both the anycast-based approach and the GCD approach. Using the MAnycastR deployment, with 32 VPs, we measure the site enumeration capabilities of both approaches when using the same VPs. We target nameservers that support CHAOS records, and measure the number of distinct CHAOS responsive values observed under the assumption that different values indicate different sites, Figure 10 plots the name server IPv4 prefixes by number of unique CHAOS records observed (x-axis) against the site counts found with the different methodologies (y-axis). We observe two things. First, for low numbers of distinct CHAOS records, both the anycast- and GCD-based approaches estimate a slightly higher number of sites. We do not have a firm explanation for this but speculate that this maybe a case of name server infrastructures where CHAOS records are used to differentiate between co-located servers at a single location (values such as ‘auth1’ and ‘auth2’ suggest this), but not between anycast sites. Second, we observe that the anycast-based approach consistently approximates the assumed CHAOS “ground truth” more closely suggesting the anycast-based approach may be slightly better at estimating a lower bound for the number of anycast sites, especially for cases where the anycast-based approach receives responses at many vantage points.
@@ Appendix D External dataset
As mentioned in §5.7, BGPTools uses the anycast-based approach to detect anycast. When an address is classified as anycast they assume the entire BGP prefix announcing this address is anycast. In this section we show this assumption does not hold by comparing both daily censuses using data from the 20th of December 2024.

### v3

@@ 8. Discussion
Lessons learned – We find the performance of both anycast-based and GCD measurements depend on the number, and geographical and topological diversity of sites. Since the majority of anycast is from hypergiants with global deployments, we observe that few nodes on different continents can detect the majority of anycast, as also exemplified by BGPTools (BGPTools, 2025). However, for detecting regional anycast a larger measurement platform is required. Furthermore, enumeration and geolocation of anycast sites requires a well-distributed measurement platform.
@@
We find our census provides good coverage of the anycast landscape, as confirmed by extensive validation. However, for IPv6 we are restricted by hitlist coverage. Additionally, for enumeration we provide a lower bound as latency-based measurements struggle to differentiate between anycast sites in near-proximity.
Whilst the ACs found using the GCDLS scans (targeting the entire hitlist) can be used solely as input for the daily GCD scans (targeting only ACs), we find the daily anycast-based measurements (targeting the entire hitlist) allows us to cover prefixes that switch from unicast to anycast. Additionally, we provide coverage of partial anycast prefixes, i.e., prefixes containing both anycast and unicast addresses, using GCDIPv4 scans.
@@
When using our census, we advise to keep the FP limitation of the anycast-based approach in mind, especially when two VPs receive responses. As for GCD, we find it is highly accurate but has rare cases of FNs due to regional anycast.
@@
We encourage operators to run their own census using the LACeS tool, which supports both anycast- and latency-based measurements. Furthermore, our daily AC lists are made available in our public census, making it possible to combine them with other AC lists to improve coverage.
Community value – Both operators, governments, and academics benefit from knowing which services are anycasted (and where they are located) to perform risk resilience assessment, determine regulatory compliance, and much more. Furthermore, by performing the census daily we track such issues longitudinally. We foresee that the dataset may have values for studies that explore, e.g., Internet resilience or locations of Internet services. We commit to continuously provide this service to the community in years to come and make the dataset publicly available.
Limitations – LACeS has limitations that affect the coverage of the daily census we provide. As we intend on providing this daily census for the foreseeable future, we address those limitations here and indicate future plans to resolve them.
LACeS only covers prefixes on hitlists containing hosts responsive to its probing methods. We intend to add probing methods and contribute to improving hitlists. For daily coverage provided by the anycast-based measurements, we require an anycast deployment with topological diversity. We are investigating the use of BGP communities and additional providers to improve coverage and reduce false positives. For GCD, accuracy is dependent on the number and distribution of VPs, as Ark grows we expect improved coverage of regional anycast. Finally, we intend on extending LACeS by performing GCD using DNS.
@@ 9. Conclusions and Future Work
We built LACeS, a new system that combines the MAnycast2 anycast measurement approach of Sommese et al. (Sommese et al., 2020) and the Great Circle Distance (GCD) approach pioneered by Cicalese et al. with iGreedy (Cicalese et al., 2015) to build an accurate and efficient daily anycast census. Through extensive validation we show that LACeS allows for an efficient, accurate and precise daily anycast census for both IPv4 and IPv6 and with support for transport layer probing using UDP/DNS and TCP. We intend to provide daily LACeS censuses to the community for the foreseeable future, covering 17+months as of Sep. ’25, through a public Git repository (git, 2025) and release the entire LACeS toolchain under a permissive open source license (Hendriks, 2025a).
We intend to further extend LACeS by including a trigger-based detection of anycast not visible with daily census granularity, e.g., using BGP route collectors, and provide an API to the community for live measurement of anycast. Finally, we are planning to use LACeS to detect suspected BGP hijacking.
Acknowledgements.
This work was supported by the RIPE NCC Community Fund, INTERSCT (NWO grant NWA.1160.18.301), the National Science Foundation (NSF grant OAC-2131987), and the European Commission’s GÉANT GN5-2 Project. We thank the operators that shared ground truth data with us, including Wouter de Vries (Cloudflare), Stephen Strowes (Fastly), Colin Petrie (Global NTT), Marco Davids (SIDN), Josh Simpson (InternetNZ), Carlos Collao Vilches (Punktum), Thomas Dupas (DNS Belgium), Jaromir Talir and Tomáš Hála (CZ.NIC). We also thank Oliver Gasser (IPInfo) for sharing their detected anycast prefixes and Thijs van den Hout (SIDN) for performing LACeS measurements using their ccTLD deployment. Finally, we thank our shepherd for their guidance and all reviewers whose feedback helped improve this paper.
Appendix A Ethics
Our ethics review board does not require a specific ethics approval for measurements conducted according to community best practices. In this work, we follow these practices by ensuring that the address space of our measurement infrastructure has correct pointers to abuse contacts and references to a page explaining the goal of the measurements, with clear instruction for opting out. The analysis conducted in this paper was also helpful in establishing a significantly lower and manageable probing rate. In the early phase of creating LACeS, we prioritized such analysis which allowed us to design LACeS to be responsible and minimize the performance impact on targeted infrastructure. Figure 4 shows these early measurements at the top of the figure.
Appendix B GCD using RIPE Atlas
@@
We performed GCD using RIPE Atlas for the 23,821 ACs of 16 September, 2024. This measurement lasted three days due to rate-limiting, even with increased spending quotas. Furthermore, it incurred a cost of 37 million RIPE Atlas credits which would take a volunteer hosting a RIPE Atlas probe approximately five years to accumulate. For these reasons we find RIPE Atlas unsuitable for the daily census.
@@
However, as discussed in §5.2 we do find that RIPE Atlas achieves better enumeration. Yet, this comes at a significant price in terms of probing cost. This can be seen in Figure 11, where we plot the number of PoPs detected for a Cloudflare prefix, with presence in 300+ cities, and the probing cost when decreasing the inter-node distance. Our RIPE Atlas scan was performed using 481 VPs with at least 100km inter-VP distances to maximize geographical coverage. By increasing the inter-VP distance up to 1,000km we reduce the number of VPs. We find that there is a linear increase in enumeration capabilities, whereas the probing cost shows an exponential increase.
@@ C.1. Detection
As mentioned in §2.3, prior work attempted to detect and enumerate anycast using CHAOS queries. We build on this prior work and leverage CHAOS queries as an opportunity to perform a side-by-side comparison in isolation of the anycast-based approach in LACeS and the built-in GCD approach. We take our nameserver hitlist as a starting point. We send CHAOS queries to each target from each of our 32 VPs according to RFC 4892 (Woolf and Conrad, 2007) to the IPv4 and IPv6 addresses on the hitlist and record the responses. Then we run a separate synchronised anycast-based measurement with 1 second offsets against the hitlist as well as a separate GCD-based measurement. In other words: we run both the anycast-based and GCD measurement toward all addresses using the LACeS deployment. Of the 161 K nameservers probed, we detect 2,762 to be anycast using the anycast-based approach of which 2,371 are also found using GCD. Counting the number of unique CHAOS records, we find 2,404 out of the 161 K nameservers return multiple distinct records. However, as noted by Fan et al. (Fan et al., 2011), these include load-balanced nameservers at a single location that return multiple records. Taking the intersection of nameservers with multiple CHAOS records and those detected as anycast using the anycast-based methodology, we find 1,082 prefixes. Conversely, 1,175 probed nameservers do not support CHAOS-class records and 147 are either falsely classified as anycast or configured with the same record replicated at all locations.
We also observe 414 nameservers that are unresponsive to ICMP and TCP, hence the GCD method cannot detect them. 50 of these prefixes are detected to be anycast using the anycast-based method, we observe they give strong indication of being TPs as they are captured at ¿ 3 VPs. These 50 include prefixes from well-known operators, such as eBay and Oracle, highlighting the utility of the added protocol support in LACeS.
@@ C.2. Enumeration
As mentioned, the LACeS tool has built-in support for both the anycast-based approach and the GCD approach. Using the LACeS deployment, with 32 VPs, we measure the site enumeration capabilities of both approaches when using the same VPs. We target nameservers that support CHAOS records, and measure the number of distinct CHAOS responsive values observed under the assumption that different values indicate different sites, Figure 12 plots the name server IPv4 prefixes by number of unique CHAOS records observed (x-axis) against the site counts found with the different methodologies (y-axis). We observe two things. First, for low numbers of distinct CHAOS records, both the anycast- and GCD-based approaches estimate a slightly higher number of sites. This is unsurprising, as previous work has shown that CHAOS records are often used to differentiate between co-located servers (Fan et al., 2011) (e.g., we observe values such as ‘auth1’ and ‘auth2’ suggesting this behavior) where they combined it with traceroute to avoid overestimation. Second, we observe that the anycast-based approach consistently approximates the assumed CHAOS “ground truth” more closely suggesting the anycast-based approach may be slightly better at estimating a lower bound for the number of anycast sites, especially for cases where the anycast-based approach receives responses at many vantage points. Despite CHAOS records being a poor indicator of anycast, we intend on including it in our daily scanning as it provides insightful information for nameservers that implement this.
@@ Appendix D External dataset
As mentioned in §5.8, BGPTools uses the anycast-based approach to detect anycast. When an address is classified as anycast they assume the entire BGP prefix announcing this address is anycast. In this section we show this assumption does not hold by comparing both daily censuses using data from the 20th of Dec. ’24.
@@
Figure 13. LACeS  detection of anycast candidates for ICMPv4, TCPv4, and DNSv4.
Figure 14. LACeS  detection of anycast candidates for ICMPv6, TCPv6, and DNSv6.
Appendix E Protocol coverage
To extend coverage compared to iGreedy and MAnycast2 we added TCP and DNS over UDP probing. Figure 13 shows a breakdown of which prefixes can be detected using our anycast-based approach with different protocols using a so-called upset plot. First, the bottom left shows the totals for each protocol, i.e., we detect 30,609 anycast prefixes with ICMP, 11,569 with TCP, and 2,997 with DNS. Next, below the histogram we indicate the intersection shown by each bar. We color this by three categories: blue for results of a single protocol, green for groupings of two protocols, and red for the intersection of all three protocols. The bars show the totals and percentage that each intersection gives, e.g., the first bar shows that 19,095 prefixes (57.7% of the total) were discovered to be anycast using only ICMP (i.e., TCP and DNS did not detect these prefixes as anycast). Next, we have the prefixes detected as anycast by both ICMP and TCP, totaling to 8,815, etc. We also include the IPv6 Upset plot in Figure 14.
Appendix F Disclaimer
Any views and opinions expressed in this work are those of the authors and do not necessarily reflect those of the European Union, the European Research Council, NSF, or the Dutch research funding agency NWO. Neither the European Union, nor NSF, nor NWO can be held responsible for them.